A forecast is to be made of the results of 5 cricket matches which can be a win, a draw , or a loss. Find the no of forecasts containing 0,1,2,3,4,5 errors respectively.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jee king

for each match we have 3 possibbilities 1 is for true forecast and 2 for wromg forecast.

0 error

all forecast must be true =1*1*1*1*1=1

1 error

total ways =5C1 *2*1*1*1*1 =10

2 error

total ways =5C2 *2*2*1*1*1 =40

3 error

total ways =5C3 *2*2*2*1*1 =80

4 error

total ways =5C4 *2*2*2*2*1 =80

5 error

total ways =5C5 *2*2*2*2*2 =32


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin